OpenStudy (anonymous):

Use the dartboard on the right (description ; a basic dartboard with a 19 for bullseye and 8 for anywhere else around). Can you score exactly 100 points with 7 darts that land on the board? Explain

OpenStudy (wolf1728):

Here's a good diagram:

OpenStudy (gorv):

let hit on bulls eye=x

OpenStudy (gorv):

so remain will be hit else where =7-x dart

OpenStudy (gorv):

total point for bulleyes=x*19=19x for other=(7-x)*8=56-8x so total=100 100=19x+56-8x subtracting 56 both side 100-56=19x-8x=11x 44=11x divide both side by 11 44/11=x 4=x yeah its possible if you hit 4 times bulls eye